# Fully Open Source Moxin-LLM Technical Report

### v1

@@ 3.2.1 Text Data
However, subsequent analyses reveal a significant limitation: some corpora within RedPajama contain a large percentage of duplicate content. The deduplication guidelines in RedPajama operate only within individual data sources, leaving inter-source duplicates largely unaddressed. To improve data quality and training efficiency, SlimPajama was developed as a refined iteration of RedPajama, offering a cleaned and extensively deduplicated version [50].
@@
Additionally, we utilize the DCLM-BASELINE dataset [38], which is derived from CommonCrawl, a web-crawled dataset [57]. The construction of DCLM-BASELINE involves several steps. First, resiliparse is employed to extract text from CommonCrawl. Second, deduplication is performed using MinHash [58] within a suffix array pipeline [59, 36] and near-duplicate Bloom filtering, which enhances the exact document and paragraph deduplication scheme [43]. Third, recent studies [60, 43, 61] demonstrate that utilizing learnable models as quality filters leads to downstream performance improvements. Consequently, DCLM-BASELINE applies a fastText OH-2.5 combined with an ELI5 classifier score to retain the top 10% of documents.
@@ 3.2.2 Coding Data
Programming is crucial for LLMs to support various downstream tasks, such as code completion from natural language descriptions, documentation generation for individual functions, and auto-completion of code snippets. Furthermore, as code is generally better structured and organized than natural language, training on code data may improve the LLM reasoning capabilities [62]. Therefore, We use part of the-stack-dedup dataset [63] during the pretraining.
@@ 4.3 Alignment Evaluation
| Moxin-7B-chat | 6.42 |
| Llama 2 13B Chat | 6.65 |

### v2

@@ 3.2.1 Text Data
However, subsequent analyses reveal a significant limitation: some corpora within RedPajama contain a large percentage of duplicate content. The deduplication guidelines in RedPajama operate only within individual data sources, leaving inter-source duplicates largely unaddressed. To improve data quality and training efficiency, SlimPajama111https://huggingface.co/datasets/cerebras/SlimPajama-627B was developed as a refined iteration of RedPajama, offering a cleaned and extensively deduplicated version [50].
@@
Additionally, we utilize the DCLM-BASELINE [38] dataset222https://huggingface.co/datasets/mlfoundations/dclm-baseline-1.0, which is derived from CommonCrawl, a web-crawled dataset [57]. The construction of DCLM-BASELINE involves several steps. First, resiliparse is employed to extract text from CommonCrawl. Second, deduplication is performed using MinHash [58] within a suffix array pipeline [59, 36] and near-duplicate Bloom filtering, which enhances the exact document and paragraph deduplication scheme [43]. Third, recent studies [60, 43, 61] demonstrate that utilizing learnable models as quality filters leads to downstream performance improvements. Consequently, DCLM-BASELINE applies a fastText OH-2.5 combined with an ELI5 classifier score to retain the top 10% of documents.
@@ 3.2.2 Coding Data
Programming is crucial for LLMs to support various downstream tasks, such as code completion from natural language descriptions, documentation generation for individual functions, and auto-completion of code snippets. Furthermore, as code is generally better structured and organized than natural language, training on code data may improve the LLM reasoning capabilities [62]. Therefore, We use part of the-stack-dedup [63] dataset333https://huggingface.co/datasets/bigcode/the-stack-dedup during the pretraining.
@@ 4.3 Alignment Evaluation
| Moxin Chat 7B | 6.42 |
| Llama 2 Chat 13B | 6.65 |
@@
4.4 Generation Examples
We demonstrate some text generation examples from our Moxin Chat 7B and Llama2 Chat 7B in Table 4.4 and 4.4.
| Example 1 | ||||||||||||||||||||||||||||||
| Moxin Chat 7B |
| Llama2 Chat 7B |
Generation Examples.
[t] Example 2 Moxin Chat 7B Can you explain the concept of normalization in machine learning? Normalization is a technique used to scale the data before using it for training a machine learning model. It is used to make the data more uniform, and to prevent it from being skewed by large or small values. There are several ways to normalize data, including min-max scaling, z-score normalization, and logarithmic scaling. In min-max scaling, the values are scaled to a range between 0 and 1, by dividing them by the maximum value in the dataset. In z-score normalization, the values are scaled to have a mean of 0 and a standard deviation of 1, by subtracting the mean from each value and dividing by the standard deviation. In logarithmic scaling, the values are transformed using the logarithm function, which can be useful for normalizing data with a wide range of values. Normalization is an important step in preparing data for machine learning, as it can help to improve the performance of the model by making the data more uniform and easier to analyze. It can also help to prevent overfitting and improve generalization. Llama2 Chat 7B Can you explain the concept of normalization in machine learning? Normalization is a technique used in machine learning to transform input data into a range of values that are suitable for use in a model. The goal of normalization is to ensure that the data is on the same scale, so that the model can learn meaningful relationships between the features. There are several types of normalization techniques, including: 1. Min-max normalization: This involves scaling the data so that it falls within a specific range, usually between 0 and 1. 2. Standardization: This involves subtracting the mean and dividing by the standard deviation of the data, so that the data has a mean of 0 and a standard deviation of 1. 3. Z-score normalization: This involves subtracting the mean and dividing by the standard deviation of the data, so that the data has a mean of 0 and a standard deviation of 1. 4. Log transformation: This involves taking the logarithm of the data, which can help to reduce the effect of extreme values in the data. 5. Square root transformation: This involves taking the square root of the data, which can help to reduce the effect of outliers in the data. Normalization is an important preprocessing step in machine learning, as it can help to improve the performance of the model by reducing the effect of irrelevant features and improving the interpretability of the results. Here are some key points to keep in mind when normalizing data for machine learning: 1. Normalization can improve the performance of the model by reducing the effect of irrelevant features and improving the interpretability of the results. 2. There are several types of normalization techniques, including min-max normalization, standardization, z-score normalization, log transformation, and square root transformation. 3. The choice of normalization technique will depend on the nature of the data and the specific requirements of the model. 4. Normalization can be applied to both continuous and categorical data. 5. It is important to normalize the data before training the model, as normalization can improve the convergence of the model and reduce the risk of overfitting. 6. Normalization can also be used to improve the interpretability of the results by reducing the effect of irrelevant features and improving the scalability of the model. 7. It is important to carefully evaluate the performance of the normalized data and consider the trade-offs between different normalization techniques. 8. Normalization can be applied to different features of the data, such as the input features, the target variable, or the output features. 9. Normalization can be used in combination with other preprocessing techniques, such as feature scaling, feature selection, and data augmentation. 10. Normalization is a powerful tool for improving the performance of machine learning models, but it is important to carefully evaluate the performance of the normalized data and consider the trade-offs between different normalization techniques. Generation Examples.